OpenStudy (anonymous):

f y = 8x − 2, which of the following sets represents possible inputs and outputs of the function, represented as ordered pairs? {(0, −2), (1, 6), (2, 14)} {(0, −2), (2, 13), (−5, −40)} {(1, 7), (2, 7), (3, 7)} {(1, 6), (2, 13), (3, 24)}

OpenStudy (anonymous):

f(x) is also y, but we should use f(x) here, because x will keep track of what we use. f(x) = 8x − 2 x = 0 8(0) = 0 - 2 = -2 the first is (0,-2) x = 1 8(1) = 8 - 2 = 6 the second one is (1, 6) could you see what it is if x = 2 to get the final answer ?

OpenStudy (anonymous):

A? @silverr

OpenStudy (godlovesme):

https://web.geogebra.org/app/# use the link ^^^

OpenStudy (anonymous):

sorry, yes, @TheHangingTree !!